# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v3

@@ V-D Compute and Cost Estimates
On 8 NVIDIA A100 GPUs, running 128 jobs in parallel, we generated 3,200 vulnerable samples in 2h57m. This averages to 3s per injection with concurrency. At Maryland’s commercial electricity rate (14.82¢/kWh), this corresponds to a cost of about \$0.44 per 1,000 injections (GPU power only). These numbers show that AVIATOR can effectively scale dataset generation at very low cost.
@@ VI Discussion
This section reflects on why AVIATOR works, its practical impact, limitations, and outlook.
@@ VI-A Why Does AVIATOR Work?
We identify the core design choices that explain AVIATOR’s effectiveness in vulnerability injection. These include its agentic decomposition, targeted model-level adaptation, iterative refinement capabilities, and integration of static analysis feedback.
(1) AVIATOR benefits from agentic decomposition. The injection task is divided among 13 specialized agents, each handling specific operations like vulnerability localization, injection, or diff-based verification. This modular approach enhances accuracy significantly, as demonstrated by improvements from 31% to 91% on FormAI and from 84% to 95% on SARD100.
(2) AVIATOR improves through targeted fine-tuning of a single critical agent. Rather than retraining all models, we fine-tune only the Vulnerability Injection agent using supervised fine-tuning (SFT) with LoRA. This lightweight adaptation requires less than 10 hours of training on a single NVIDIA A100 GPU and improves injection accuracy from 85% to 91% on FormAI. The model is trained on PrimeVul, a dataset entirely disjoint from the test sets, showing that this adaptation generalizes well across domains.
(3) AVIATOR leverages iterative workflow refinement to improve injection robustness. Agents operate in feedback loops to enhance injection robustness. For example, the Diff Agent triggers refinement cycles when necessary, boosting accuracy up to 10 points on FormAI and 4 points on SARD100. Notably, AVIATOR achieves a strong baseline even without fine-tuning, provided the underlying LLM has sufficient coverage of the target language, highlighting that its workflow design is the primary driver of performance. This makes the framework readily portable to new programming languages and under-studied vulnerability classes, while ensuring diversity in the types and structures of injected flaws.
(4) AVIATOR integrates traditional static analysis tools to validate injection quality. The Verifier Agent uses cppcheck to detect whether the injected code violates real-world safety constraints. Its feedback is interpreted and used by subsequent agents to iteratively refine the injection.
@@ VI-B Limitations and Threats to Validity
Our evaluation focuses on C/C++ using three datasets (SARD100, FormAI, and PrimeVul), which may not capture all real-world diversity. To mitigate this, we combined synthetic and real-world benchmarks of varying complexity.
@@ VII CONCLUSION
Our experimental evaluations, conducted across diverse benchmarks, demonstrate that AVIATOR seems to significantly outperform existing techniques, achieving high success rates in vulnerability injection. The modular design of AVIATOR, combined with cutting-edge techniques such as Retrieval-Augmented Generation and Low-Rank Adaptation, enables the creation of large-scale datasets. These datasets are not only accurately labeled but are diverse, making them very useful for training robust and effective vulnerability detection models.
The contributions of this work advances the state-of-the-art in automated vulnerability injection, while providing a scalable solution to the pressing challenge of data scarcity in software security research.
Future work will explore further enhancements in agentic workflows and fine-tuning strategies, as well as further evaluation on real-world code, to broaden the scope and applicability of AVIATOR in practical scenarios.
VIII Data Availability
The datasets and code used in this study will be made publicly available upon acceptance in a curated replication package. This package will include: (i) the implementation of the AVIATOR workflow, (ii) the prompts used for each agent in the workflow, (iii) the benchmarks used in our experiments with references to their original sources, and (iv) the processed splits and retrieval knowledge bases used in AVIATOR. All materials will be released under an open-source license.

### v4

@@ V-E Compute and Cost Estimates
On 8 NVIDIA A100 GPUs, running 128 jobs in parallel, we generated 5,000 vulnerable samples in 4h37m. This averages to 3s per injection with concurrency. At Maryland’s commercial electricity rate (14.82¢/kWh), this corresponds to a cost of about \$0.44 per 1,000 injections (GPU power only). These numbers show that AVIATOR can effectively scale dataset generation at very low cost.
V-F RQ5. Impact on Downstream Vulnerability Detection
To assess AVIATOR’s practical utility, we evaluate how augmenting training data with AVIATOR-generated vulnerable samples impacts downstream deep learning-based vulnerability detection (DLVD) performance. Table VII reports the performance metrics for four distinct deep learning architectures (Devign, Reveal, LineVul, CodeT5) trained on the Devign and PrimeVul datasets augmented with generated samples using different strategies.
Consistent downstream gains and comparison to augmentation baselines
[…]
TABLE VI: Overall Average Relative Improvement of AVIATOR Compared to Other Strategies.
Compared Strategy Precision Recall F1-Score
No Aug +12.2% +104.9% +21.8%
ROS +9.0% +26.0% +12.5%
VulGen +30.3% +5.7% +25.2%
VGX +39.2% -25.4% +24.6%
VulScribeR +1.4% +7.4% +2.9%
TABLE VII: Downstream vulnerability detection performance (Precision, Recall, F1). Each model is trained on the seed dataset augmented with 5k injected vulnerable samples using each strategy. Best F1 scores per model and setting are highlighted in bold. Results for non-AVIATOR strategies are taken from the VulScribeR paper [10]; we report new results for AVIATOR and CodeT5.
Strategy Devign Model Reveal Model LineVul Model CodeT5 Model
P R F1 P R F1 P R F1 P R F1
Seed Dataset: Devign; Test Dataset: BigVul Test
No Aug 6.27 43.24 10.95 7.28 77.27 13.31 7.90 29.19 12.43 6.08 65.71 11.13
ROS 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93 7.52 51.81 13.13
VulGen 6.49 62.32 11.75 6.36 70.59 11.66 8.51 22.15 12.29 7.34 31.05 11.87
VGX 6.06 59.30 10.99 6.89 73.29 12.60 5.35 98.82 10.15 7.20 80.17 13.21
VulScribeR 8.79 29.41 13.53 11.64 38.00 17.82 11.43 18.81 14.22 7.09 76.09 12.97
AVIATOR 9.59 24.80 13.83 9.19 56.94 15.82 11.27 29.99 16.38 8.08 65.89 14.39
Seed Dataset: PrimeVul Train; Test Dataset: PrimeVul Test
No Aug 25.93 5.12 8.55 7.35 44.63 12.62 18.31 31.15 23.06 19.34 25.68 22.07
ROS 7.45 52.44 13.04 7.58 43.66 13.07 18.48 30.05 22.89 18.03 28.42 22.07
VulGen 4.94 59.76 9.11 5.29 59.27 9.72 19.39 28.23 22.99 19.82 24.04 21.73
VGX 5.61 53.66 10.15 5.44 50.73 9.82 19.77 27.69 23.06 16.61 32.79 22.05
VulScribeR 7.57 55.37 13.31 7.81 49.76 13.51 17.29 31.88 22.42 20.26 25.14 22.44
AVIATOR 7.73 51.82 13.45 8.65 36.32 13.97 19.74 28.05 23.18 16.96 31.33 22.01
AVIATOR leads to an average F1 score improvement of +21.8% relative to the original dataset (No Aug), confirming that the injected samples provide substantial additional training signal beyond the original data. Compared to the Random Oversampling baseline (ROS), AVIATOR achieves a +12.5% average F1 score improvement. Notably, ROS itself outperforms VulGen and VGX by 9.0% and 8.5% in average F1, respectively, highlighting the strength of AVIATOR beyond simple class-balancing effects.
When compared to pattern-based injection methods, AVIATOR improves average F1 score by +25.2% over VulGen and by +24.6% over VGX. Among all baselines, VGX is the only method that achieves a higher average recall than AVIATOR; however, this comes at the cost of a substantial drop in precision. This behavior is consistent with VGX’s reliance on single-statement injection templates and its limited coverage of vulnerability classes (23 CWEs). Such augmentation strategies can cause detectors to become overly sensitive (high recall) while substantially increasing false positives (low precision) which ultimately limits their practical utility for large-scale data augmentation [10]. In contrast, AVIATOR improves F1 through a more balanced precision-recall profile, indicating that it provides a cleaner and less artifact-driven training signal.
Finally, AVIATOR outperforms VulScribeR by +2.9% in average F1, with gains driven primarily by a substantial improvement in recall (+7.4%) while maintaining, and even slightly improving, precision (+1.4%). This precision-recall behavior indicates that AVIATOR’s augmentation introduces less label noise and achieves broader coverage of vulnerable patterns than VulScribeR. In particular, AVIATOR provides detectors with more learnable and discriminative vulnerable signals, enabling them to identify a wider set of true vulnerabilities (reflected by higher recall) without incurring additional false alarms (as evidenced by stable precision). Given VulScribeR’s reported injection success rate of approximately 82%, the observed precision-recall profile is consistent with AVIATOR generating higher-fidelity vulnerable samples on average, with an estimated success rate of 93%, thereby strengthening downstream generalization across detector families.
Overall, these findings suggest that the data generated by AVIATOR is of consistently higher quality than existing augmentation baselines, as reflected by its stable and robust downstream performance across models and evaluation settings.
Distributional fidelity and realism
To further assess the realism of AVIATOR-generated data and explain its downstream gains, we analyzed the distributional properties of the augmented datasets. Figure 5 reports the Kullback-Leibler (KL) divergence between the original training data and the augmented data, using two views: sample counts and entropy distributions across CWE categories. A lower KL divergence indicates that the augmentation strategy preserves the natural shape and diversity distribution of real-world datasets, rather than introducing artificial distortion.
AVIATOR achieves the lowest KL divergence under both metrics, which confirms that it remains closest to the original distribution. VulGen and VGX introduce notably larger shifts, consistent with augmentation that concentrates on a narrower set of injection patterns. VulScribeR reduces this distortion compared to pattern-mining baselines but still deviates more than AVIATOR. Overall, these distributional results align with the precision-recall trends in Table VI: lower distortion corresponds to more realistic, learnable injected vulnerabilities that improve generalization without inducing excessive false alarms.
Refer to caption Refer to caption
(a) CWE Entropy Distribution Shift to the (b) CWE Sample Count Distribution
Original Dataset (x-axis truncated). Shift to the Original Dataset.
Figure 5: Distributional analysis of augmented datasets. KL divergence measures the deviation from the original, natural distribution of vulnerabilities (lower is better). AVIATOR introduces the least distributional distortion.
Operational superiority: quality, cost, and efficiency
Beyond downstream metrics, AVIATOR demonstrates significant operational advantages over the state-of-the-art VulScribeR in terms of data quality and cost-efficiency.
  • High Syntactic Validity: We utilized the fuzzy C parser Joern [51] to verify the validity of the augmented samples, following the methodology in [10]. While VulScribeR reports a filtration rate (samples rejected due to syntax errors) of 2%–13%, AVIATOR achieves a significantly lower rejection rate: 1.13% for Devign and 1.98% for PrimeVul. This confirms that AVIATOR’s agentic workflow produces more usable training samples with less formatting and parsing noise.
[…]
@@ VI Discussion
This section reflects on why AVIATOR is effective, what its results imply for vulnerability dataset construction and downstream detection, and the main limitations of the current study.
@@ VI-A Why Does AVIATOR Work?
We identify the core design choices that explain AVIATOR’s effectiveness in vulnerability injection. These include its agentic decomposition, targeted model-level adaptation and iterative refinement capabilities, mirroring how human experts inject vulnerabilities in practice.
(1) Agentic decomposition provides reliable guidance and controllability. AVIATOR decomposes vulnerability injection into 13 specialized agents, each responsible for a well-defined subtask such as code understanding, localization, transformation, diff-based checking, and refinement. This explicit separation enables structured reasoning and targeted self-correction instead of relying on a single monolithic generation step. Empirically, the workflow design drives most of the gains: injection success on FormAI increases from 31% (single-step baseline) to 91%, and on SARD100 from 84% to 95%. These results indicate that decomposition and tool-augmented verification improve injection fidelity.
(2) Targeted fine-tuning strengthens the core transformation step. Instead of retraining the full pipeline, we fine-tune only the Vulnerability Injector with LoRA-based SFT. This lightweight adaptation (under 10 hours on a single A100 GPU) improves FormAI from 85% to 91% while reducing variance. The model is trained on PrimeVul, a dataset entirely disjoint from the test sets, showing that this adaptation generalizes well across domains.
(3) AVIATOR leverages iterative workflow refinement to improve injection robustness. Agents operate in feedback loops to enhance injection robustness. AVIATOR explicitly detects and corrects common failure modes in LLM code editing (e.g., cosmetic edits, truncation, or insufficient modifications). For example, the Diff Agent can trigger additional refinement passes when changes are overly disruptive or insufficient. This mechanism improves robustness by avoiding brittle, single-statement edits and contributes to consistent performance even without fine-tuning, provided the base LLM has reasonable language coverage. As a result, AVIATOR can adapt to new vulnerability classes and languages, while maintaining diversity in injected flaws.
(4) High-quality injection translates into downstream generalization. KL-divergence analysis shows that AVIATOR induces the smallest shift in CWE sample-count and entropy distributions relative to the original training set. In addition to the high injection success rate, this result supports the downstream precision-recall profile: methods that over-concentrate augmentation on narrow templates tend to distort the training distribution and encourage detectors to produce too many false positives. AVIATOR instead adds signal while tracking the baseline distribution more closely, which plausibly explains its stronger and more stable F1 score gains across detector families.
(5) Operational scalability strengthens practical impact. AVIATOR maintains a low syntax rejection rate under Joern-based filtering ($2\%$) while reducing generation cost by $4.3\times$ compared to VulScribeR. These properties enable large-scale dataset construction and augmentation without accumulating unusable samples or incurring prohibitive cost, and they make AVIATOR suitable for repeated regeneration under consistent quality controls.
@@ VI-B Limitations and Threats to Validity
Our evaluation focuses on C/C++ using three datasets (SARD100, FormAI, and PrimeVul), which may not capture all real-world diversity. To mitigate this, we combined synthetic and real-world benchmarks of varying complexity. Extending AVIATOR to multi-file contexts and additional languages remains future work.
Function-Level Context Limitations
Our current instantiation operates and validates injections at the function level. For datasets such as PrimeVul, missing project context (e.g., type definitions, globals, external callees) can make some injected samples not determinable under function-local analysis. Accordingly, our manual evaluation assesses whether the intended weakness is present within the function body and marks cases that depend on unavailable context as unknown, as done in prior studies [40]. Ongoing work extends AVIATOR toward repository-level injection and validation.
@@
Distribution metrics and interpretation
Entropy- and KL-based analyses capture distributional distortion and code diversity proxies, not semantic equivalence to real vulnerabilities. These metrics support our downstream conclusions, but they do not replace direct correctness validation. We therefore interpret them as complementary evidence rather than as ground-truth realism measures.
Cost estimates and hardware dependence
Our cost comparison focuses on generation-time compute under a specific hardware and parallelization setting. Absolute costs vary with deployment, batching, and model choice. The key point is the relative trend: AVIATOR achieves strong quality while enabling lower-cost scaling than large proprietary API-based LLM augmentation pipelines.
@@ VII CONCLUSION
Across three widely used benchmarks (SARD100, FormAI, and PrimeVul), AVIATOR achieves high injection success rate (93% across benchmarks), outperforming prior injection techniques and providing stronger guarantees for constructing benchmark datasets with reliable labels. Beyond injection correctness, AVIATOR also delivers the strongest downstream benefits under the standard DLVD augmentation protocol: injecting 5k samples improves average downstream F1 score by +21.8% over no augmentation and +24.6% over VGX, while surpassing the prior downstream state of the art VulScribeR by +2.9% average F1 and +7.4% recall without degrading precision. AVIATOR induces the lowest distributional distortion relative to the original training dataset, supporting the conclusion that it generates more realistic and lower-noise vulnerable samples that generalize across detector families.
[…]
Overall, these results show that AVIATOR advances the state of the art in automated vulnerability injection by jointly improving injection fidelity, downstream utility, and scalability. Future work will extend AVIATOR beyond function-level edits toward multi-function and project-level injection, enabling validation under richer context and supporting a broader class of real-world vulnerability scenarios.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
Data will be made available on request.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author(s) used Gemini in order to order to receive suggestions on wording and form of certain paragraphs. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
CRediT authorship contribution statement
Amine Lbath: Writing - Original Draft, Writing – review & editing, Conceptualization, Methodology, Software, Data Curation, Investigation, Visualization. Massih-Reza Amini: Writing - Original Draft, Writing – review & editing, Conceptualization, Methodology, Supervision. Aurelien Delaitre: Writing – review & editing, Conceptualization, Methodology, Supervision, Validation. Vadim Okun: Writing – review & editing, Conceptualization, Methodology, Supervision, Validation.